# A Review of Bayesian Uncertainty Quantification in Deep Probabilistic Image Segmentation

### v5

@@ 7.5 Segmentation backbone
Due to emphasis on the uncertainty methods, the backbone feature extractor is often an overlooked element in uncertainty modeling. For example, CNN-based encoder-decoder models such as the U-Net, remain the preferred backbone architecture (Eisenmann et al., 2023). Nevertheless, Vision Transformers (ViTs) (Dosovitskiy et al., 2020) such as SegFormer, Swin U-Net, Mask2Former already outperform CNNs in general segmentation problems (Xie et al., 2021; Zhang & Yang, 2021; Cheng et al., 2022). With the exception of experiments in Kahl et al. (2024) (combining SSN (Monteiro et al., 2020) with HRNet (Yuan et al., 2020b)), all employed models use a CNN backbone. The limited adoption of ViTs likely stems from CNNs’ advantageous inductive biases, whereas Transformers typically require extensive pretraining on large datasets (Caron et al., 2021). Therefore, we recommend integrating the discussed methods into similar architectures to achieve both strong classification performance and effective ambiguity modeling. Also, CNNs have benefited from the recent developments in Transformers. For instance, ConvNext takes inspiration from contemporary Transformers to modernize existing ResNet-based CNNs, retaining the inductive biases of convolutional filters and achieving significant performance gains (Liu et al., 2022c). These same inductive biases have been added to ViTs through CNN adapters to also enhance performance Chen et al. (2023b).
7.6 Modeling feature distributions
In many instances, latent-variable models are employed to model the observer variability. To aid with appropriate model selection, we will discuss the strengths and weaknesses of each of those approaches. A summary of this discussion is presented in Table 7. Note that, while classified PixelCNNs and SSNs are classified as pixel-level methods, both can also be phrased as a latent-variable model.
ELBO suboptimality
Known for its flexibility to various datasets, fast sampling time and the interpretable latent space, VAEs seem to be the most popular choice for aleatoric uncertainty quantification. Nonetheless, the shortcomings of VAEs are well known. For example, such models suffer from inference suboptimality related to ELBO optimization Cremer (2018); Zhao et al. (2017) and literature on the VAE-based PU-Net often describe behavior similar to the well-known phenomena of model collapse Valiuddin et al. (2024a; b); Qiu & Lui (2020). This has been hypothesized to be caused by excessively strong decoders Chen et al. (2016) and is especially apparent when dealing with complex hierarchical decoding structures, where additional modifications such as the GECO objective Kohl et al. (2019a), residual connections Kohl et al. (2019a); Baumgartner et al. (2019), or deep supervision Baumgartner et al. (2019) are required for generalization. A unique benefit of this approach is the ability to semantically interpret the latent space with, for example, interpolation between annotation styles or the exploration of low-likelihood regions.
Limitation of sequential modeling
The adoption of sequential sampling models DDPMs or PixelSeg is rather limited compared to the VAE-based models. This is unexpected, as both models outperforms the latter. This is likely due to their tedious sequential inference procedure Song et al. (2020); Zheng et al. (2023); Meng et al. (2023); Zhang et al. (2022c), a crucial limitation that is exacerbated in supervised settings, which often require validation through sampling on a separate data split. Despite these shortcomings, the strengths of DDPMs should not be overlooked. Their iterative sampling allows for highly flexible modeling and preserves high-frequency details often lost in latent-variable approaches like VAEs, which can produce blurry reconstructions. Furthermore, recent advances in faster sampling methods, such as DDIMs (Song et al., 2020) and Flow Matching Lipman et al. (2022), make these models promising candidates for further exploration.
Discrete vs. continuous
It can be noted in Table 6 that the best-performing DDPM models are discrete in nature Chen et al. (2023a); Zbinden et al. (2023). While it is debatable whether shifting to categorical distributions is required for complex image generation Chen et al. (2022). Visually, this can be already quite apparent for the multi-class case, where the transition to noise is visually more gradual in the discrete transition (see Figure 13). However, Table 1 shows AB Chen et al. (2022) performs almost identically to CCDM Zbinden et al. (2023), suggesting that straightforward thresholding of continuous models have little-to-no performance sacrifice compared with explicitly modeling discrete distributions. These observation call for further investigation of the merits of categorical distributions in segmentation settings.
(a) LIDC-IDRI Armato (2011).
(b) CityScapes Cordts et al. (2016a).
Figure 13: Continuous vs. categorical forward diffusion process with cosine noise scheduling Nichol & Dhariwal (2021). Note the use of categorical diffusion that results in a more gradual transition for the multi-class case.
Others
Stochastic Segmentation Networks (SSNs) are a simple, fast, and model-agnostic approach for modeling uncertainty. However, they can suffer from training instability due to the requirement that the covariance matrix be invertible. To address this limitation, the authors propose masking out the background to avoid exploding variances and using uncorrelated Gaussians when the covariance matrix is singular. Despite these efforts, SSNs scale poorly with data, and a potential solution is to shift the stochasticity deeper into the network by introducing a latent prior, resembling VAE-based models but with a more complex latent space. CAR, on the other hand, is a less commonly used method, likely due to the well-known instability of GAN-based models Arjovsky & Bottou (2017). GANs often require additional heuristic terms in the loss function to ensure stable training. CAR is no exception to this, as its objective function combines four separate losses.
Recommendation
We advocate for a deeper contextualization of contemporary generative modeling research within probabilistic segmentation models. Generative modeling is a rapidly evolving field and any such model can be modified for segmentation through intricate conditioning. For example, DDPMs were initially proposed for general image generation (Ho et al., 2020), but was quickly adopted for stochastic segmentation. Furthermore, flow-based models also enjoy many benefits yet to be explored in this field, such as exact likelihood modeling and faster sampling Lipman et al. (2022). In terms of model selection, we have noticed SSN being the most practical solution across literature due to being model agnostic and receiving good benchmarking scores (Kahl et al., 2024; Monteiro et al., 2020; Ng et al., 2022).
Table 7: Comparison between generative models that model aleatoric uncertainty.
Method Advantages Disadvantages Examples
SSN Model agnostic, explicit likelihoods, fast sampling possibly unstable, memory intensive (Monteiro et al., 2020; Kahl et al., 2024; Zepf et al., 2023a; Philps et al., 2024; Kahl et al., 2024)
PixelCNN Explicit, exact likelihoods Sequential sampling, memory intensive Zhang et al. (2022c)
GAN Fast sampling, flexible Unstable training, poorly defined objective, implicit likelihoods (Kassapis et al., 2021; Isola et al., 2017)
VAE Fast sampling, flexible, interpretable latent space, ELBO Mode/posterior collapse, amortization gap (Kohl et al., 2019b; Selvan et al., 2020; Valiuddin et al., 2021; 2024a; Qiu & Lui, 2020; Bhat et al., 2022a; 2023; Valiuddin et al., 2024b; Zhang et al., 2022b; Kohl et al., 2019a; Baumgartner et al., 2019; Hu et al., 2023; Liu et al., 2022b; Fischer et al., 2023; Savadikar et al., 2021; Rafael-Palou et al., 2021; Viviers et al., 2023b; a; Chotzoglou & Kainz, 2019; Saha et al., 2021a; b; Long et al., 2021a; Schmidt et al., 2023; Savadikar et al., 2021; Rafael-Palou et al., 2021; Zepf et al., 2024; 2023a; Hu et al., 2022)
DDPM Flexible, expressive Sequential sampling (Chen et al., 2023a; Zbinden et al., 2023; Rahman et al., 2023)
7.7 Modeling parameter distributions
In the previous sections, we reviewed feature modeling methods, typically used to quantify observer variability. Other downstream applications usually rely on epistemic uncertainty modeling, i.e. modeling parameter distributions, which will be discussed in this section.
Criticism of MC Dropout
Nearly all literature on epistemic uncertainty quantification relies on approximations of variational inference (VI). Among these, MC Dropout is the most commonly used due to its simplicity, low computational cost, and ease of implementation. However, MC Dropout has faced substantial criticism Folgoc et al. (2021); Osband (2016); Kingma et al. (2015); Gal et al. (2017a). For instance, it can assign zero probability to the true posterior or introduce erroneous multi-modality Folgoc et al. (2021). It is also sensitive to model size and dropout rate rather than the observed data Osband (2016); Kingma et al. (2015). To address these issues, some alternatives attempt to learn the dropout rate Kingma et al. (2015); Gal et al. (2017a). Still, uncertainty estimates from MC Dropout should be treated as a convenient byproduct—not a reliable core feature. When precise uncertainty quantification is critical, such methods are generally insufficient and best avoided. When uncertainty modeling is used to boost model performance, rigorous benchmarking of uncertainty often becomes secondary. In such cases, methods like MC Dropout are treated as regularization tools being part of the deep learning toolbox, rather than for genuine uncertainty estimation. However, it’s debatable whether the performance gains stem from better uncertainty modeling or from other factors. For instance, both ensembling and MC Dropout introduce a model combination effect—effectively averaging multiple networks. MC Dropout also resembles placing an $L_{2}$ penalty on weights. While these effects can improve generalization, similar gains might be achieved with simpler, more efficient regularizers. Perhaps it is the stochasticity itself that helps. Still, if performance is the main goal, uncertainty modeling is likely not the most efficient route due to its added computational cost. Instead, improved performance may simply be a side effect of these techniques.
Comparative studies
Finally, the optimal method for uncertainty quantification has yet to be determined. In some works, MC dropout was found to perform better than ensembling Hoebel et al. (2020); Roy et al. (2018b), while in other works ensembling excels Mehrtash et al. (2020); Ng et al. (2022). Finally, there is convincing evidence to prefer Concrete Dropout rather than MC Dropout Mukhoti & Gal (2018). All things considered, the preference for a particular methodology seems to carry a strong data-dependency Dahal et al. (2020); Jungo & Reyes (2019). A valuable contribution to the field pertains a comprehensive benchmark paper, which compares all available epistemic uncertainty quantification methods across a wide range of datasets. In particular, such study can elucidate the data-dependent preference for specific methodologies (i.e. why ensembling or MC Dropout is often preferred). Additionally, recent studies have shown the benefits of moving from a few large, to many small experts when using an MoE ensemble for language modeling (He, 2024). Future work should also experiment with this.
Alternatives to Variational Inference
@@ 7.7 Modeling parameter distributions
Approaches besides VI, such as Markov Chain Monte Carlo (MCMC) or Laplace Approximations, are also viable options to approximate the Bayesian posterior. Especially the Laplace Approximation can be very beneficial, as it is easily applicable to pretrained networks. Notably, both Laplace Approximation and VI are biased and operate in the neighborhood of a single mode, while MCMC methods are a more attractive option when expecting to fit multi-modal parameter distributions. Zepf et al. (2023b) explored the impact of the Laplace Approximation on segmentation networks; however, beyond this work, the method has received limited attention in the field.
@@
The multiple forward passes required in many Bayesian uncertainty quantification can incur cumbersome additional costs. Hence, considerable efforts have been made towards sampling-free uncertainty models (Mukhoti et al., 2023; Liu et al., 2020; Van Amersfoort et al., 2020; Postels et al., 2019), only depending on a single forward pass. Mukhoti et al. (2023) show that Gaussian Discriminant Analysis after training with SoftMax predictive distribution can in some instances surpass methods such as MC Dropout and ensembling. Along similar lines, Evidential Deep Learning also possesses the advantage of quantifying both uncertainties with a single forward pass. This framework is based on a generalization of Bayes theorem, known as the Dempster-Shafer Theory of Evidence (DST) (Dempster, 1967). While common in Bayesian probability, DST does not require prior probabilities and bases subjective probabilities on belief masses assigned on a frame of discernment, i.e. the set of all possible outcomes. The use of Evidential Deep Learning has seen success in conventional classification problems (Sensoy et al., 2018), and Ancha et al. (2024) recently applied this concept to segmentation to decouple aleatoric and epistemic uncertainty within a single model. However, follow-up research in this direction remains limited to date.

### v8

@@ 7.7 Segmentation backbone
Due to emphasis on the method, task and application, the backbone feature extractor is often an overlooked element in uncertainty modeling. For example, CNN-based encoder-decoder models such as the U-Net, remain the preferred backbone architecture (Eisenmann et al., 2023). Nevertheless, Vision Transformers (ViTs) (Dosovitskiy et al., 2020) such as SegFormer, Swin U-Net, Mask2Former already outperform CNNs in general segmentation problems (Xie et al., 2021; Zhang and Yang, 2021; Cheng et al., 2022). With the exception of experiments in Kahl et al. (2024) (combining SSN (Monteiro et al., 2020) with HRNet (Yuan et al., 2020b)), all employed models use a CNN backbone. The limited adoption of ViTs likely stems from CNNs’ advantageous inductive biases, whereas Transformers typically require extensive pretraining on large datasets (Caron et al., 2021). Nonetheless, CNNs have benefited from the recent developments in Transformers. For instance, ConvNext takes inspiration from contemporary Transformers to modernize existing ResNet-based CNNs, retaining the inductive biases of convolutional filters and achieving significant performance gains (Liu et al., 2022c). These same inductive biases have been added to ViTs through CNN adapters to also enhance performance (Chen et al., 2023b). Building upon these observations, we conclude that future research should actively explore the integration of state-of-the-art Vision Transformer (ViT) and hybrid CNN-Transformer backbones to advance the field.
7.8 Contemporary uncertainty
While current efforts frequently leverage established uncertainty modeling techniques for segmentation, the practical integration involves significant challenges related to ensuring architectural and task compatibility. To accelerate progress, the segmentation community should actively explore and integrate newly developed uncertainty quantification methodologies, rather than relying solely on standard, established practices.
VI alternatives
@@ 7.8 Contemporary uncertainty
Approaches besides VI, such as Markov Chain Monte Carlo (MCMC) or Laplace Approximations, are also viable options to approximate the Bayesian posterior. Especially the Laplace Approximation can be very beneficial, as it is easily applicable to pretrained networks. Notably, both Laplace Approximation and VI are biased and operate in the neighborhood of a single mode, while MCMC methods are a more attractive option when expecting to fit multi-modal parameter distributions. Zepf et al. (2023b) explored the impact of the Laplace Approximation on segmentation networks. However, beyond this work, the method has received limited attention in the field.
@@
The multiple forward passes required in many Bayesian uncertainty quantification can incur cumbersome additional costs. Hence, considerable efforts have been made towards sampling-free uncertainty models (Mukhoti et al., 2023; Liu et al., 2020; Van Amersfoort et al., 2020; Postels et al., 2019), only depending on a single forward pass. Mukhoti et al. (2023) show that Gaussian Discriminant Analysis after training with a SoftMax predictive distribution can in some instances surpass methods such as MC Dropout and ensembling. Along similar lines, Evidential Deep Learning also possesses the advantage of quantifying both uncertainties with a single forward pass. This framework is based on a generalization of Bayes theorem, known as the Dempster-Shafer Theory of Evidence (DST) (Dempster, 1967). While common in Bayesian probability, DST does not require prior probabilities and bases subjective probabilities on belief masses assigned on a frame of discernment, i.e. the set of all possible outcomes. The use of Evidential Deep Learning has seen success in conventional classification problems (Sensoy et al., 2018), and Ancha et al. (2024) recently applied this concept to segmentation attempt to decouple aleatoric and epistemic uncertainty within a single model. However, follow-up research in this direction remains limited to date.